How can i complete the square with :
2k^2 + 4k-10=0
my answer was 5, it's obviously wrong... help plz! thank you

Answer by nerdybill(7384)

2k^2 + 4k-10=0
divide through by 2:
k^2 + 2k-5=0
k^2 + 2k = 5
take 1/2 'b', square it and add to both sides:
k^2 + 2k + 1 = 5 + 1
k^2 + 2k + 1 = 6
(k+1)(k+1) = 6
(k+1)^2 = 6
k+1 = sqrt(6)
k = +-sqrt(6) - 1
k = {-sqrt(6)-1, sqrt(6)-1}